# EchoGuide: Active Acoustic Guidance for LLM-Based Eating Event Analysis from Egocentric Videos

### v1

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 International Symposium on Wearable Computers; October 8–12, 2024; Melbourne, Australiabooktitle: Proceedings of the 2024 International Symposium on Wearable Computers (ISWC ’24), October 5–9, 2024, Melbourne, Australiaprice: 15.00doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/XX/$X$
@@ 1. Introduction
Smart glasses have emerged as a promising wearable form factor for tracking eating behaviors. To alleviate the need for manual recording, various sensing systems based on smart glasses have been developed to distinguish eating behavior from arm movements (Zhang et al., 2022) or facial muscle movements(Shin et al., 2022). However, most of these systems can only identify when eating occurs but not what is being eaten, which is critical information for interpreting eating behaviors. Conversely, sensing systems such as cameras, which can capture detailed information on eating (e.g., the type of food consumed), have high power consumption, making continuous operation impractical on commodity smart glasses.
@@
Active acoustic sensing(Li et al., 2022; Zhang et al., 2023; Mahmud et al., 2023, 2024) has been shown as a new expressive and efficient modality for detecting and recognizing full-body motion using wearable devices, especially in low-power and low-compute settings. The latest work ActSonic(Mahmud et al., 2024) has shown that using active-acoustic sensing on glasses can recognize over 28 types of everyday activities (including eating) in the wild with 89% F1 score at each second without the need for any training data from a new user. More specifically, it recognizes eating activities with an F1 score of 90% in completely unconstrained environments. However, this sensing modality doesn’t capture the full context of a given activity. For activity recording and downstream applications (such as calorie counting or recipe assistance), it’s important to understand not only what action (e.g., when eating happens) is performed given body motion, but also what objects (e.g., what is the food being eaten) the action is being performed with.
@@ 2. Related Work
2.1. Multimodal Image/Video Captioning and Summarization
As larger-scale language and multimodal generative “foundation models” (Bommasani et al., 2021) have been trained and released, image and video captioning has extended from simply determining the similarity of images/videos to a premade list of captions (Wray et al., 2016, 2019; Wang et al., 2022; Radford et al., 2021; Ni et al., 2022; Lin et al., 2022) towards generating captions for new videos based on either fine-tuning inexpensive smaller-scale Large Language Models with video encoders and captions (Zhao et al., 2023) or leveraging the emergent properties of natively multimodal Large Language models (such as the OpenAI GPT-4 multimodal series (OpenAI et al., 2024; GPT, [n. d.]) and Google Gemini multimodal series (Team et al., 2023)), truly “open-world” video captioning becomes more possible especially in a “zero-shot” paradigm without labeled examples. These systems can be applied offline over the course of a video to create “activity records”: long documents which encode which activities a person might be completing within the course of a video, and which can be efficiently indexed and searched.
Extracting insights from preprocessed ”activity records” requires methods which can generate relevant answers to queries that are grounded in specific documents. Recent generative methods, especially in scenarios involving domain-specific information, leverage the Retrieval-Augmented Generation (Lewis et al., 2020) for returning helpful responses given queries and documents containing relevant information. These methods primarily involve chunking and embedding input documents into a searchable index leveraging large pre-trained neural embedding models, searching and returning the content of the top ”k” chunks based on embedding/semantic similarity with input questions, and leveraging the returned content as part of a generated response.
@@ 2.1. Multimodal Image/Video Captioning and Summarization
2.2. Eating Recognition on Glasses
Various sensing modalities on eyeglasses form factors have been proposed to track eating events. These modalities include EMG electrodes (Zhang and Amft, 2017), piezoelectric sensing (Shin et al., 2022; Farooq and Sazonov, 2016), contact microphones (Bi et al., 2018), microphones and IMUs (Mirtchouk et al., 2017), and sensor fusion (Bedri et al., 2020, 2022). While these systems track eating episodes, they lack the ability to provide detailed information related to eating activities, such as what food a person eats. This limitation is due to the absence of optimized access to an egocentric camera for extended monitoring periods.
@@ 3.2.1. Using Active Acoustic Sensing to localize relevant actions and clip videos
We leverage ActSonic’s ResNet-18 model as an event detector by splitting the active acoustic differential echo profile(Mahmud et al., 2024) (synchronized with the video) into consecutive 2-second windows which can be passed directly into the model. We define a set of “domain-specific classes” within the label space of ActSonic which capture important events for this particular domain, extract class predictions for all sliding windows (essentially treating the class prediction’s “timestamp” as the last timestamp of the corresponding window), and construct intervals of events by filtering for “domain-specific classes” and joining equivalent predictions in consecutive windows to create clips without requiring dense captions.
@@ 4. User Study
4.1. Participants
The EchoGuide user study received approval from the Institutional Review Board for Human Participant Research (IRB) at our organization. We recruited 10 participants for a semi-in-the-wild user study at their homes. However, 1 participant’s data was lost during the user study. Therefore, we ended up with 9 valid participants in the study, ranging in age from 19 to 34. 6 participants self-reported as male while 3 self-reported as female. We collected basic demographic data and their ratings on the prototype through an IRB-approved questionnaire. The average comfort rating on a Likert scale of 0 to 5 was 2.62.
4.2. Data Capture Apparatus
We captured acoustic data using the sensing system integrated into EchoGuide eyeglasses and recorded egocentric activity video data via the EchoGuide GoPro Hero9 (GoPro, 2020) camera mounted on the participants’ heads using a lightweight body mount from the same manufacturer. The camera’s horizontal and vertical field of view was set to $118\degree118$ ° and $69\degree69$ ° respectively. It recorded egocentric videos at a resolution of 720720720720p and a frame rate of $3030$ fps.
4.3. Study Procedure
We conducted a 9-participant semi-in-the-wild user study in unconstrained environments such as participants’ homes and offices. The recruited participants were equipped with eyeglasses and a head-mounted camera. We synchronized the acoustic and video data with a clap action. After synchronization, the participants were allowed to continue with their daily routines without any interruption. Upon completion of the 40-minute study, the participants returned the devices to the researcher. In total, we had 5 participants collect data at their homes and 4 participants in their office environments.
@@ 5.1.1. Answer alignment with dense captioning
Given a single question related to the domain, we find semantic similarity between the answer from a RAG QA agent that has indexed an activity record with alternative sampling (e.g. leveraging the ultrasonic modality to filter and caption fewer frames) and the answer from a RAG QA agent that has indexed an activity record with dense video sampling (e.g. captioning the entire video, which can reduce overall efficiency). Semantic similarity is captured via BERT F-1 scores (Zhang* et al., 2020), which captures pairwise cosine similarities between BERT output embeddings to capture semantic and contextual information and which shows high correlation with human evaluations on summarization and captioning tasks (closely related to this work).
BERT F-1 score operates within a cosine similarity range of -1 to 1, with higher numbers considered as more similar. As an example, given a “ground-truth” densely-captioned document’s answer of “What did C eat or drink in this session?” as “Person C poured and drank tea and ate food with chopsticks.”, sentences which contain as much relevant information from the input sentence while excluding information unseen in the input, such as “Person C was seen pouring tea into a cup, holding a cup of tea, and scooping food with chopsticks”, will have higher BERT F-1 scores (due to including specific details such as drinking tea or using chopsticks) when compared with “Person C ate and drank a variety of items throughout the day”.
@@ 5.3. Quantitative Results
Table 2. Results of manual evaluation of EchoGuide + GPT4o given 1fps vs 0.5 fps sampling of frames from ActSonic-defined clips. Notation is defined as (X/Y) where X=accuracy at 1fps and Y=accuracy at 0.5fps
@@ 6.1. Metrics/Evaluation Procedure
EchoGuide, however, focuses not only on providing general summaries via activity records of an individual’s day from video and wearable sensor data, but also on answering targeted questions about these summaries by leveraging the image-text pertaining of large multimodal language models. We evaluate this method via manual review and annotation of the system’s answers to three eating questions (“What did C eat/drink? What utensils did C use while eating/drinking? What container did C eat or drink out of?”) when configured to use GPT4o (GPT, [n. d.]) to caption images sampled at two varying FPS levels (1fps and 0.5fps) from clips proposed by ActSonic, and report accuracy metrics showing whether EchoGuide extracts correct values for these questions as compared to manually-determined ”ground truth” (taken by watching the reference video and determining which item is present): we’ve shown accuracy values given 1fps and 0.5fps sampling in Table 2. Accuracy values are defined as a 0/1 binary: 0 represents responses that do not have complete overlap with the ground-truth, while 1 represents responses that do have a complete overlap with the ground truth.
@@ 7. Discussion
7.1. Further Reduction on video recording when deployed in the wild
Our study results showed that EchoGuide helped reduce video recordings by an average of 68% without significantly impacting the quality of summarization for eating activities. However, we want to point out that this percentage of reduction will likely be significantly higher if the system is deployed for full-day recording. In the user study, we only asked participants to collect data for 40 minutes, including the meal. Consequently, the ratio of eating activities in our dataset is significantly higher than it would be in a full day of recording. Therefore, if ActSonic is used to only activate the camera during eating activities in a full-day recording, the data reduction will likely be significantly higher than 68%. Additionally, the frame rate of recording can be further reduced to answer specific questions, saving energy and processing resources. We plan to explore these questions further in the next step.
7.2. Comparison to Egocentric Video recorded on glasses
@@
The initial hardware system for EchoGuide was not collected using “camera-enabled smartglasses” such as the Meta Ray-Bans (Meta, 2023) due to limitations on recording videos with these off-the-shelf smart glasses. Instead, we used a head-mounted GoPro to easily capture egocentric activity videos. We found the information related to eating captured by smart glasses and our GoPro settings to be highly similar. To help readers understand the difference in images captured by these two devices, we used RayBan Smart glasses and a GoPro HERO9 mounted on the head to capture the same dietary scenario (a drink on the table), as shown in Figure 3. We found that the GoPro has a much wider field of view and can capture more general scene details in the orientation used by EchoGuide compared to the camera on the Meta RayBan smart glasses. However, because most foods are present near the center of the field of view, the difference in view angle between the two cameras did not impact the captured information. Therefore, the results reported in the paper can still be referenced for egocentric video analysis captured on smart glasses.
7.3. Exploring additional domains for EchoGuide
While EchoGuide was adapted to focus primarily on localizing and understanding eating activities from video and acoustics, we also run a separate exploratory study with three participants operating in both the “eating” and “cooking” domains. Each participant engaged in 5 sessions of 4 minutes each within each domain with interventions between sessions to stop and restart data collection, for a total of 40 minutes per participant.
@@ 7.3. Exploring additional domains for EchoGuide
7.4. Improving comfort and reliability of hardware prototype
The current hardware prototype leverages a Teensy-based microcontroller situated on the left side of the eyeglasses which is connected to a phone for power and recording control, along with a Go-Pro head-mounted camera for video capture. Leveraging a lower-power BLE with a LiPo battery (similar to Google Glass) as the primary microcontroller module for active acoustic recording, along with a Flex-PCB that reduces extraneous wiring, can reduce the unwieldy nature of the acoustic system. Developing custom low-power, high-FPS and high-resolution cameras (such as event-based cameras) that are purpose-built for eyeglass frames can also enable seamless video recording without a GoPro requirement, reducing the weight of the systems considerably. Building eyeglass frames that can swap lenses in a custom way, or building a system that can be seamlessly applied on any eyeglass, can reduce the likelihood of participants with custom prescriptions being unable to see through the provided eyeglasses.
7.5. Reducing software latency to enable real-time applications
Currently, EchoGuide processes and asks questions over activity records in an offline fashion, but many users may want to understand their activities in an online fashion (for instance, asking about previous meals while evaluating what food to get at a restaurant). As seen in other concurrent works (Zhang et al., 2023), active acoustic postprocessing could be completed on a smartphone, and with advancements in embedded AI chips and stronger networking modules for more robust cloud access, it may be possible to do end-to-end inference online with both edge-deployed and cloud-deployed models.
7.6. Improving overall model flexibility to new situations
While Sec 5 and Table 1 show promising results for EchoGuide usage (combining video and ultrasonic modalities) across two distinct domains and procedural styles in everyday activities, further improvements can be made to enhance overall system performance. Collecting and fine-tuning on a larger base dataset of ultrasonic captures of activities can enable more robust, user-independent detection of human body motion, while leveraging steadily more powerful large multimodal models can enable more robust and generalizable video captions that encode more domain-specific or estoeric information.

### v2

@@ Abstract.
Eating Detection; Acoustic Sensing; Activity Recognition; Foundation Models
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM International Symposium on Wearable Computers; October 5–9, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 2024 ACM International Symposium on Wearable Computers (ISWC ’24), October 5–9, 2024, Melbourne, VIC, Australiadoi: 10.1145/3675095.3676611isbn: 979-8-4007-1059-9/24/10ccs: Computing methodologies Machine learningccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
Smart glasses have emerged as a promising wearable form factor for tracking eating behaviors. To alleviate the need for manual recording, various sensing systems based on smart glasses have been developed to distinguish eating behavior from arm movements (Zhang et al., 2022), ambient sound(Thomaz et al., 2015) or facial muscle movements(Shin et al., 2022). However, most of these systems can only identify when eating occurs but not what is being eaten, which is critical information for interpreting eating behaviors. Conversely, sensing systems such as cameras, which can capture detailed information on eating (e.g., the type of food consumed), have high power consumption, making continuous operation impractical on commodity smart glasses.
@@
Active acoustic sensing(Wang et al., 2018) has been shown as a low-power and powerful sensing modality for tracking and interpret various types of fine-grained body poses on wearables, including facial expressions(Li et al., 2022), gaze(Li et al., 2024b), finger pose(Yu et al., 2024; Lee et al., 2024), body pose(Mahmud et al., 2023), tongue gesture(Sun et al., 2023), silent speech recognition(Zhang et al., 2023b, a), authentication(Li et al., 2024a; Chen et al., 2024b) and physiological signal (Chen et al., 2024a; Fan et al., 2023). The latest work ActSonic(Mahmud et al., 2024) has shown that using active-acoustic sensing on glasses can recognize over 28 types of everyday activities (including eating) in the wild with 89% F1 score at each second without the need for any training data from a new user. More specifically, it recognizes eating activities with an F1 score of 90% in completely unconstrained environments. However, this sensing modality doesn’t capture the full context of a given activity. For activity recording and downstream applications (such as calorie counting or recipe assistance), it’s important to understand not only what action (e.g., when eating happens) is performed given body motion, but also what objects (e.g., what is the food being eaten) the action is being performed with.
@@ 2. Related Work
Multimodal Image/Video Captioning and Summarization: As larger-scale language and multimodal generative “foundation models” (Bommasani et al., 2021) have been trained and released, image and video captioning has extended from simply determining the similarity of images/videos to a premade list of captions (Wray et al., 2016, 2019; Wang et al., 2022; Radford et al., 2021; Ni et al., 2022; Lin et al., 2022) towards generating captions for new videos based on either fine-tuning inexpensive smaller-scale Large Language Models with video encoders and captions (Zhao et al., 2023) or leveraging the emergent properties of natively multimodal Large Language models (such as the OpenAI GPT-4 multimodal series (OpenAI et al., 2024; GPT, [n. d.]) and Google Gemini multimodal series (Team et al., 2023)), truly “open-world” video captioning becomes more possible especially in a “zero-shot” paradigm without labeled examples. These systems can be applied offline throughout a video to create “activity records”: long documents which encode which activities a person might be completing within the course of a video, and which can be efficiently indexed and searched.
Extracting insights from preprocessed ”activity records” requires methods which can generate relevant answers to queries that are grounded in specific documents. Recent generative methods, especially in scenarios involving domain-specific information, leverage the Retrieval-Augmented Generation (Lewis et al., 2020) for returning helpful responses given queries and documents containing relevant information.
@@
Eating Recognition on Glasses: Various sensing modalities on eyeglasses form factors have been proposed to track eating events. These modalities include EMG electrodes (Zhang and Amft, 2017), piezoelectric sensing (Shin et al., 2022; Farooq and Sazonov, 2016), contact microphones (Bi et al., 2018), microphones and IMUs (Mirtchouk et al., 2017), and sensor fusion (Bedri et al., 2020, 2022). While these systems track eating episodes, they lack the ability to provide detailed information related to eating activities, such as what food a person eats. This limitation is due to the absence of optimized access to an egocentric camera for extended monitoring periods.
@@ 3.2.1. Using Active Acoustic Sensing to localize relevant actions and clip videos
We leverage ActSonic’s ResNet-18 model as an event detector by splitting the active acoustic differential echo profile (synchronized with the video) into consecutive 2-second windows which can be passed directly into the model. We define a set of “domain-specific classes” within the label space of ActSonic which capture important events for this particular domain, extract class predictions for all sliding windows (essentially treating the class prediction’s “timestamp” as the last timestamp of the corresponding window), and construct intervals of events by filtering for “domain-specific classes” and joining equivalent predictions in consecutive windows to create clips without requiring dense captions.
@@ 4. User Study
Participants The EchoGuide user study received approval from the Institutional Review Board for Human Participant Research (IRB) at our organization. We recruited 10 participants for a semi-in-the-wild user study at their homes. However, 1 participant’s data was lost during the user study. Therefore, we ended up with 9 valid participants in the study, ranging in age from 19 to 34. 6 participants self-reported as male while 3 self-reported as female. We collected basic demographic data and their ratings on the prototype through an IRB-approved questionnaire. The average comfort rating on a Likert scale of 0 to 5 was 2.62.
Data Capture Apparatus We captured acoustic data using the sensing system integrated into EchoGuide eyeglasses and recorded egocentric activity video data via the EchoGuide GoPro Hero9 (GoPro, 2020) camera mounted on the participants’ heads using a lightweight body mount from the same manufacturer. The camera’s horizontal and vertical field of view was set to $118\degree118$ ° and $69\degree69$ ° respectively. It recorded egocentric videos at a resolution of 720720720720p and a frame rate of $3030$ fps.
Study Procedure We conducted a 9-participant semi-in-the-wild user study in unconstrained environments such as participants’ homes and offices. The recruited participants were equipped with eyeglasses and a head-mounted camera. We synchronized the acoustic and video data with a clap action performed by the researcher as the two sensors were physically separated. After synchronization, the participants could continue normal activities without interruption if they ate or drank at least one item within the 40-minute window. Upon completing the 40-minute study, the participants returned the devices to the researcher. We had 5 participants collect data at their homes and 4 in their office environments.
@@ 5.1.1. Answer alignment with dense captioning
Given a single question related to the domain, we find semantic similarity between the answer from a RAG QA agent that has indexed an activity record with alternative sampling (e.g. leveraging the ultrasonic modality to filter and caption fewer frames) and the answer from a RAG QA agent that has indexed an activity record with dense video sampling (e.g. captioning the entire video, which can reduce overall efficiency but captures all possible information). Semantic similarity is captured via BERT F-1 scores (Zhang* et al., 2020), which captures pairwise cosine similarities (within the range -1 to 1) between BERT output embeddings to capture semantic and contextual information and which shows high correlation with human evaluations on summarization and captioning tasks (closely related to this work). Similarity scores are used to quantify information loss between the densely captioned model and the ActSonic-captioned model.
@@ 5.3. Quantitative Results
Table 2. Results of manual evaluation of EchoGuide + GPT4o given 1fps vs 0.5 fps sampling of frames from ActSonic-defined clips (based on zero-shot accuracy). Notation is defined as (X/Y) where X=accuracy at 1fps and Y=accuracy at 0.5fps
@@ 6.1. Metrics/Evaluation Procedure
EchoGuide, however, focuses not only on providing general summaries via activity records of an individual’s day from video and wearable sensor data, but also on answering targeted questions about these summaries by leveraging the image-text pertaining of large multimodal language models. We evaluate this method via manual review and annotation of the system’s answers to three eating questions (“What did C eat/drink? What utensils did C use while eating/drinking? What container did C eat or drink out of?”) when configured to use GPT4o (GPT, [n. d.]) to caption images sampled at two varying FPS levels (1fps and 0.5fps) from clips proposed by ActSonic, and report accuracy metrics showing whether EchoGuide extracts correct values for these questions as compared to manually-determined ”ground truth” (taken by watching the reference video and determining which item is present): we’ve shown accuracy values given 1fps and 0.5fps sampling in Table 2. Accuracy values are defined as a 0/1 binary: 0 represents responses that do not overlap with the ground truth, while 1 represents responses that do completely overlap with the ground truth.
@@ 7. Discussion
Further Reduction on video recording when deployed in the wild Our study results showed that EchoGuide helped reduce video recordings by an average of 68% without significantly impacting the quality of summarization for eating activities. However, we want to point out that this percentage of reduction will likely be significantly higher if the system is deployed for full-day recording. In the user study, we only asked participants to collect data for 40 minutes, including the meal. Consequently, the ratio of eating activities in our dataset is significantly higher than it would be in a full day of recording. Therefore, if ActSonic is used to only activate the camera during eating activities in a full-day recording, the data reduction will likely be significantly higher than 68%. Additionally, the frame rate of recording can be further reduced to answer specific questions, saving energy and processing resources. We plan to explore these questions further in future works
@@
Comparison to Egocentric Video recorded on glasses The initial hardware system for EchoGuide was not collected using “camera-enabled smartglasses” such as the Meta Ray-Bans (Meta, 2023) due to limitations on recording videos with these off-the-shelf smart glasses. Instead, we used a head-mounted GoPro to easily capture egocentric activity videos. We found the information related to eating captured by smart glasses and our GoPro settings to be highly similar. To help readers understand the difference in images captured by these two devices, we used RayBan Smart glasses and a GoPro HERO9 mounted on the head to capture the same dietary scenario (a drink on the table), as shown in Figure 3. We found that the GoPro has a much wider field of view and can capture more general scene details in the orientation used by EchoGuide compared to the camera on the Meta RayBan smart glasses. However, because most foods are present near the center of the field of view, the difference in view angle between the two cameras did not impact the captured information. Therefore, the results reported in the paper can still be referenced for egocentric video analysis captured on smart glasses.
Exploring additional domains for EchoGuide While EchoGuide was adapted to focus primarily on localizing and understanding eating activities from video and acoustics, we also run a separate exploratory study with three participants operating in both the “eating” and “cooking” domains. Each participant engaged in 5 sessions of 4 minutes each within each domain with interventions between sessions to stop and restart data collection, for a total of 40 minutes per participant.
@@
Improving comfort and reliability of hardware prototype The current hardware prototype leverages a Teensy-based microcontroller on the left side of the eyeglasses which is connected to a phone for power and recording control, along with a Go-Pro head-mounted camera for video capture. The relative weight and complexity of the combined devices were cited in user surveys as the primary reason for the low comfort rating of the prototype (as it weighed more on the head and ears).
Leveraging a lower-power BLE with a LiPo battery (similar to Google Glass) as the primary microcontroller module for active acoustic recording, along with a Flex-PCB that reduces extraneous wiring, can reduce the unwieldy nature of the acoustic system. Developing custom low-power, high-FPS and high-resolution cameras (such as event-based cameras) that are purpose-built for eyeglass frames can also enable seamless video recording without a GoPro requirement, reducing the weight of the systems considerably. Building eyeglass frames that can swap lenses in a custom way, or building a system that can be seamlessly applied on any eyeglass, can reduce the likelihood of participants with custom prescriptions being unable to see through the provided eyeglasses.
Reducing software latency to enable real-time applications Currently, EchoGuide processes and asks questions over activity records in an offline fashion, but many users may want to understand their activities in an online fashion (for instance, asking about previous meals while evaluating what food to get at a restaurant). As seen in other concurrent works (Zhang et al., 2023b), active acoustic postprocessing could be completed on a smartphone, and with advancements in embedded AI chips and stronger networking modules for more robust cloud access, it may be possible to do end-to-end inference online with both edge-deployed and cloud-deployed models.
Improving overall model flexibility to new situations While Sec 5 and Table 1 show promising results for EchoGuide usage (combining video and ultrasonic modalities) across two distinct domains and procedural styles in everyday activities, further improvements can be made to enhance overall system performance. Collecting and fine-tuning on a larger base dataset of ultrasonic captures of activities can enable more robust, user-independent detection of human body motion, while leveraging steadily more powerful large multimodal models can enable more robust and generalizable video captions that encode more domain-specific or estoeric information.
@@ 8. Conclusion
Acknowledgement
This project was supported by the National Science Foundation Grant No. 2239569 and partially by the Cornell University IGNITE Innovation Acceleration Program. We also acknowledge and thank Prof. Bharath Hariharan at Cornell CS for help with computing resources.